# Quantum amplitude estimation from classical signal processing

### v1

@@ Abstract
We demonstrate that the problem of amplitude estimation, a core subroutine used in many quantum algorithms, can be mapped directly to a problem in signal processing called direction of arrival (DOA) estimation. The DOA task is to determine the direction of arrival of an incoming wave with the fewest possible measurements. The connection between amplitude estimation and DOA allows us to make use of the vast amount of signal processing algorithms to post-process the measurements of the Grover iterator at predefined depths. Using an off-the-shelf DOA algorithm called ESPRIT together with a compressed-sensing based sampling approach, we create a phase-estimation free, parallel quantum amplitude estimation (QAE) algorithm with a total query complexity of $\sim 4.9/\varepsilon$ and a parallel query complexity of $\sim 0.40/\varepsilon$ at 95% confidence. This performance is a factor of $\times1$ and 14×14$\times$14 × improvement over Rall and Fuller [Quantum 7, 937 (2023)], for worst-case complexity, which to our knowledge is the best published result for amplitude estimation. The approach presented here provides a simple, robust, parallel method to performing QAE, with many possible avenues for improvement borrowing ideas from the wealth of literature in classical signal processing.
@@ I Introduction
Here, we demonstrate a non-iterative AE algorithm with the benefit that the number of iterations is known from the outset, so every sample can be done in parallel. In addition, the classical post-processing is a) robust to noise, allowing one to take very few samples to achieve low overall query complexity, and b) is classically efficient, not affecting the overall algorithm complexity up to log factors. Finally, in comparison to other similarly performant QAE algorithms, our approach does not require quantum signal processing (QSP) which simplifies the fault-tolerant circuits required to implemnt it. To achieve these results, we draw from ideas in (classical) signal processing, in particular from algorithms used in determining the Direction Of Arrival (DOA) of an incoming signal.
@@ II Classical signal processing based AE
When measuring |$\ket{\phi_{n}}$ in the $X$-basis (apply the Hadamard gate and then measure), we obtain |$\ket{0}$ or |$\ket{1}$ with probabilities
$p_{0}^{X}(n)$ $:=\frac{1}{2}(1+\sin((2n+1)2\theta))$ (4a)
$p_{1}^{X}(n)$ $:=\frac{1}{2}(1-\sin((2n+1)2\theta)).$ (4b)
This allows us to construct the estimator
$\frac{p_{0}^{X}(n)-p_{1}^{X}(n)}{p_{0}(n)-p_{1}(n)}=\tan((2n+1)2\theta).$ (5)
Taking the arctan of the expression above we obtain an approximation of $(2n+1)2\theta$ from which we can form the exponential
$y_{n}=e^{i(2n+1)2\theta}=e^{in4\theta+2i\theta}.$ (6)
The need to take the arctan instead of just the arccos or arcsin of Eq. (4) is needed to span the entire range of $20\leq\theta\leq\pi/$ 2.
The estimation of $y_{n}$ is not exact. The probabilities that we estimate from Eqs. (3) and (4) have binomial sampling noise. The goal of AE is to estimate $\theta$ using as few measurements as possible. In the next section, we describe a classical signal processing approach that uses the estimator defined in Eq. (6), mimicking almost exactly the same type of estimator one obtains using physical sensors that are attempting to estimate the direction of arrival of an unknown signal using noisy measurements.
@@ II.1 Subspace Based Direction Of Arrival Estimation
where we have identified a unit distance $x_{n}=(n-1)$ and angle $\omega=2\pi d\sin{\bar{\omega}}/\lambda$ for an incoming signal with wavelength $\lambda$. The parameter $y_{n}$ here is distinct from that defined in Eq. (6), but we purposely use the same notation as it turns out they are equivalent from the perspective of DOA estimation. The parameter $\epsilon_{n}$ denotes an error term due to imperfect measurements. The goal of DOA estimation, is to determine the incident angle of the incoming signal $\bar{\omega}$ using the fewest possible sensors. Looking at Eq. (6), we can view it as an incoming signal with angle of incidence $\omega=4\theta$ at position $n$. This is the main connection between AE and DOA estimation that we will use. The main difference is how we obtain noisy measurements of $y_{n}$.
@@ II.2 ESPRIT Approach to DOA Estimation
For AE, we don’t need to do the full singular value decomposition of the matrix $\boldsymbol{R}$ because we know from the outset that there is just a single incoming signal. Instead, we just need the top two eigenvectors to form the matrices $\boldsymbol{S}_{1}$ and $\boldsymbol{S}_{2}$. Because the matrix is Toeplitz, we can use the Lanczos algorithm [29] together with Fourier-transform based algorithms for computing matrix-vector products of Toeplitz matrices to obtain these two eigenvectors in time $O(M\log(M))$ [30]. Computating the pseudo-inverse of $\boldsymbol{S}_{1}$ is likewise fast, taking time $\mathcal{O}(M)$, since the matrix is of size $(M-1)\times 2$. Therefore the classical complexity of using ESPRIT for AE is $\mathcal{O}(M\log(M))$, logarithmically worse than the lower-bound query complexity of AE, but insignificant in practice. We find that the post-processing for the largest matrix we numerically simulate in this manuscript with $M=215177$ and targets a 95%-percentile error rate of $9.85\times 10^{-6}$, takes 8 seconds on a laptop with the current implementation and probably can be optimized much further.
@@ II.3 Physical and virtual arrays
Taking samples at depths of powers of two is reminiscent of the Maximum Likelihood Amplitude Estimation (MLAE) approach to AE [11]. The main difference here is the classical post-processing, which provides far better results than MLAE. This was well-known for some time in the signal processing literature [21]. We used powers of two to determine scaling factors of our approach in Sec. IV.1, but note that there is a great deal of flexibility here depending on the use-case. We highlight some alternative schedules when discussing the impact of circuit noise in Sec. IV.3.
@@ III The AE algorithm
In classical signal processing terms, we can obtain an estimate of the signal $y_{n}=e^{in4\theta+2i\theta}$ by placing a physical sensor at location $n$. Quantumly, we can obtain a estimate of the signal $y_{n}$ by taking repeated measurements of |$\ket{\phi_{n}}$ in the $Z$ and $X$ basis. As far as the DOA algorithms are concerned it does not matter how the measurements are obtained.
The next step is to determine at what depths we take measurements. The DOA algorithms expects measurements on a uniform linear array of length $M$, that is measurements of $y_{n}$ for $n\in[M]$. One may obtain this by measuring |$\ket{\phi_{n}}$ at depths all $n\in[M]$, with a query complexity of $\sum_{n=1}^{M}n=O(M^{2})$. This does not provide us with the expected quantum scaling $O(1/M)$ of the error, but rather the scaling is $O(1/\sqrt{M})$ in this case. We avoid this problem by taking samples from a far smaller set of depths and use the virtual array concept to obtain (virtual) measurements at all depths $n\in[M]$.
@@
2:  Measure the state |$\ket{\phi_{n}}$ for $n\in D$ in the $Z$ and $X$ basis and form the “signal” vector $y_{n}=e^{i\omega_{n}}$ where $\omega_{n}$ is the arctan of the expression in (5) and where the $p_{i}(n)$ and $p_{i}^{X}(n)$ are replaced by their empirical values (from the measurements).
3:  Form the signal vector $\boldsymbol{y}\in\mathbb{C}^{M}$ by computing the signal values at the virtual locations corresponding to positive virtual positions.
4:  Use the ESPRIT (or any classical DOA) algorithm with $\boldsymbol{y}$ as input to extract the angle $\hat{\omega}$.
@@ IV Numerics
Figure 3: Estimation error as a function of query complexity (top) and parallel complexity (bottom) for amplitude $a=0.5$. For csAE we simulated 500 runs giving a range of error rates (given by the vertical dots). The green triangles denote the 95-th percentile of the errors and the green fit line is used to compute the scaling factor. For chebAE we also simulated 500 runs, for a target $\varepsilon$ given by the results from csAE. The cloud of points from chebAE arise because of the iterative nature. The red and cyan stars and fit lines denote the average and worst case complexity respectively for chebAE for the 95-th percentile.
Refer to captionRefer to caption
Figure 4: Scaling factor $C_{\textrm{opt}}$ as a function of amplitude for total (top) and parallel complexity (bottom) of csAE (green triangles) and chebAE (red stars) at $\delta=0.95$ confidence. The horizontal lines correspond to the worst scaling factor for the amplitudes tested and is the number we report as the overall scaling factor.
@@
The query complexity is the number of times we use the operator $U$ (see  (1)) divided by two. Since the Grover operator $G$ consists of two calls to $U$, the query complexity will be the number of times we use $G$ with an additional number of oracle calls to $U$ at depth 0 divided by two. In chebAE [15] the number of oracle queries is defined as the number of times we apply the Grover operator $G$ ignoring the calls to $U$ at depth 0. This won’t affect constant factors in any meaningful way (we are penalizing our algorithm slightly), but the way we account for the queries seems to be fairer. We compare the two approaches ability to achieve a target $\varepsilon$ and $\delta$ on two different metrics: 1) query complexity and 2) parallel query complexity. By parallel query complexity we mean the maximum depth oracle query.
To get a better intuition of the performance of the algorithm, in Fig. 3, we provide results when estimating the amplitude $a=0.5$ for 500 random trials where the only source of randomization is the binomial sampling error when estimating the probability. We use a physical array with parameters $N_{1}=N_{2}=\cdots N_{2q}=2$ as defined in Eq. (12) for $q\in\{3,4,\ldots,8\}$. For csAE, the input is the physical sampling array, denoting the fixed query depth, as well as the number of samples to take at each depth. This results in a deterministic schedule and yields the vertical colored dots representing the estimation errors of the algorithm from different runs. The green triangles represent the $\delta=0.95$ confidence level of those runs for different depth circuits. The green line is a fit line to these 95% confidence levels.
@@
where $n_{\mathrm{max}}$ is the maximum query depth, $n$ is the current query depth, and $K$ is a constant that we chose to be $K=1.3$ for $\delta\leq 0.95$ and $K=1.8$ for $\delta=0.99$. This was determined heuristically, so it is possible that there exist better schedules $N_{\mathrm{shots}}(n)$. We note that it is also possible to have a non-deterministic schedule that utilizes information from the the top-two eigenvalues of the matrix $\boldsymbol{P}$ of Algorithm 1. When these eigenvalues are close to the remaining eigenvalues the signal and noise subspaces are not well separated. We believe that using this information could further improve the constant factor scaling, but at the cost of reducing parallelizability. We leave this to future work.
@@ IV.1 Fits
Table 1: Fitting constants for various confidence levels, along with the associated standard error obtained from the weighted-least-squares fit. Overall, csAE is comparable to the average case complexity of chebAE in terms of total query complexity, but tends to outperform in terms of worst-case complexity. In addition, csAE is highly parallelizable resulting in vast out-performance for parallel query complexity.
csAE $8.5\pm 1.6$ $0.6\pm 0.1$ $4.9\pm 0.6$ $0.40\pm 0.05$ $2.0\pm 0.1$ $0.162\pm 0.008$
chebAE (worst) $8.2\pm 0.5$ $8.0\pm 0.3$ $5.7\pm 0.2$ $5.4\pm 0.1$ $2.52\pm 0.08$ $2.23\pm 0.09$
chebAE (ave) $5.8\pm 0.4$ $5.6\pm 0.3$ $4.0\pm 0.1$ $3.77\pm 0.09$ $1.55\pm 0.08$ $1.44\pm 0.05$
@@
The performance of the algorithm depends on the value of the amplitude $a$ to be estimated. Since $a$ is an unknown parameter, we compare the performance of csAE for amplitudes in the range $[0.1,0.2,\ldots,0.9]$. For each amplitude, we perform the same analysis shown in Fig. 3 and compute the estimation error $\varepsilon_{\delta}$ at a given confidence-level $\delta$. We compute the weighted-least squares fit using both average and worse-case query complexities for each amplitude to compute the scaling factor $C_{\textrm{opt}}$.
We plot the results of these fits for $\delta=0.95$ confidence in Fig. 4 for both total query complexity and parallel query complexity using the worst-case query complexity. We report the constant factor as the maximum constant factor as a function of amplitude. We find that csAE generally performs better for larger amplitudes, while chebAE has worst-case complexity at $a=0.5$ as previously reported [15].
In Tab. 1 we provide a summary of the constant factor scaling comparing csAE to chebAE for different confidence levels and for both average and worst-case query complexity for chebAE. Overall, we find that our algorithm slightly under-performs chebAE for the average case, but over-performs for the worst case, however many of these are within the error bars.
@@
While csAE does not significantly lower the constant factor in terms of query complexity compared to chebAE, the implementation is much simpler. As with the IQAE implementation [14], the csAE approach makes use of the standard Grover iterator and do not require QSP. This implies that the circuit-level implementation is substantially simpler than for chebAE, not requiring controlled phase rotations and circuit synthesis, which adds a logarithmic factor to the required fault-tolerant circuit depth [33]. Therefore, our approach yields the best of both worlds: simple circuit construction, together with state-of-the-art query complexity. As we discuss next, our approach also offers a highly parallel implementation, something not available to either chebAE or IQAE.
IV.2 Parallelizability
Figure 5: Scaling of the constant factor for the parallel query complexity as we increase the number of quantum computers. This is for the amplitude $0.5$ and the depth schedule is $(2^{j})_{j\in[2q]}$ for $q=5$ and shots schedule $[15,13,12,11,10,8,7,6,4,3,2]$. This gives a total query complexity of $6,417$ and results in an error rate of $5.6\times 10^{-4}$ (with $95\%$ confidence). Because shots at lower $q$ values require fewer oracles calls, they can be parallelized across fewer computers resulting in the saturation near 12 QPUs.

### v2

@@ Abstract
We demonstrate that the problem of amplitude estimation, a core subroutine used in many quantum algorithms, can be mapped directly to a problem in signal processing called direction of arrival (DOA) estimation. The DOA task is to determine the direction of arrival of an incoming wave with the fewest possible measurements. The connection between amplitude estimation and DOA allows us to make use of the vast amount of signal processing algorithms to post-process the measurements of the Grover iterator at predefined depths. Using an off-the-shelf DOA algorithm called ESPRIT together with a compressed-sensing based sampling approach, we create a phase-estimation free, parallel quantum amplitude estimation (QAE) algorithm with a worst-case sequential query complexity of $\sim 4.3/\varepsilon$ and a parallel query complexity of $\sim 0.26/\varepsilon$ at 95% confidence. This performance is statistically equivalent and a 16×16$\times$16 × improvement over Rall and Fuller [Quantum 7, 937 (2023)], for sequential and parallel query complexity respectively, which to our knowledge is the best published result for amplitude estimation. The approach presented here provides a simple, robust, parallel method to performing QAE, with many possible avenues for improvement borrowing ideas from the wealth of literature in classical signal processing.
@@ I Introduction
Here, we demonstrate a non-iterative AE algorithm with the benefit that the number of iterations is known from the outset, so every sample can be done in parallel. In addition, the classical post-processing is a) robust to noise, allowing one to take very few samples to achieve low overall query complexity, and b) is classically efficient, not affecting the overall algorithm complexity up to log factors. Finally, in comparison to other similarly performant QAE algorithms, our approach does not require quantum signal processing (QSP) which simplifies the fault-tolerant circuits required to implement it. To achieve these results, we draw from ideas in (classical) signal processing, in particular from algorithms used in determining the Direction Of Arrival (DOA) of an incoming signal.
@@ II Classical signal processing based AE
Using the double angle formula
$p_{0}(n)=\frac{1+\cos((2n+1)2\theta)}{2},$ (4)
we can obtain an estimate for $\cos((2n+1)2\theta)$ by taking repeated samples of the quantum state |$\ket{\phi_{n}}$.
To apply signal processing techniques, we need to form the complex exponentials
$y_{n}=e^{i[2\theta\pi(2n+1)]}.$ (5)
Using the double angle formula, we have access to estimates of the real part of $y_{n}$, but not the imaginary part. Hence, both $y_{n}$ and $y_{n}^{*}$ are consistent with the probabilities obtained from measuring the quantum state |$\ket{\phi_{n}}$. We will show in Section III.1 how to distinguish between the two cases using classical post-processing. For now, we assume we have access to estimates of both the real and imaginary parts of $y_{n}$.
The estimation of $y_{n}$ is not exact. The probabilities that we estimate from Eqs. (3) have binomial sampling noise. The goal of AE is to estimate $\theta$ using as few measurements as possible. In the next section, we describe a classical signal processing approach that uses estimates to Eq. (5), mimicking almost exactly the same type of estimator one obtains using physical sensors that are attempting to estimate the direction of arrival of an unknown signal using noisy measurements.
@@ II.1 Subspace Based Direction Of Arrival Estimation
where we have identified a unit distance $x_{n}=(n-1)$ and angle $\omega=2\pi d\sin{\bar{\omega}}/\lambda$ for an incoming signal with wavelength $\lambda$. The parameter $y_{n}$ here is distinct from that defined in Eq. (5), but we purposefully use the same notation as it turns out they are equivalent from the perspective of DOA estimation. The parameter $\epsilon_{n}$ denotes an error term due to imperfect measurements. The goal of DOA estimation, is to determine the incident angle of the incoming signal $\bar{\omega}$ using the fewest possible sensors. Looking at Eq. (5), we can view it as an incoming signal with angle of incidence $\omega=4\theta$ at position $n$. This is the main connection between AE and DOA estimation that we will use. The main difference is how we obtain noisy measurements of $y_{n}$.
@@ II.2 ESPRIT Approach to DOA Estimation
For AE, we don’t need to do the full singular value decomposition of the matrix $\boldsymbol{R}$ because we know from the outset that there is just a single incoming signal. Instead, we just need the top two eigenvectors to form the matrices $\boldsymbol{S}_{1}$ and $\boldsymbol{S}_{2}$. Because the matrix is Toeplitz, we can use the Lanczos algorithm [29] together with Fourier-transform based algorithms for computing matrix-vector products of Toeplitz matrices to obtain these two eigenvectors in time $O(M\log(M))$ [30]. Computing the pseudo-inverse of $\boldsymbol{S}_{1}$ is likewise fast, taking time $\mathcal{O}(M)$, since the matrix is of size $(M-1)\times 2$. Therefore the classical complexity of using ESPRIT for AE is $\mathcal{O}(M\log(M))$, logarithmically worse than the lower-bound query complexity of AE, but insignificant in practice.
@@ II.3 Physical and virtual arrays
Taking samples at depths of powers of two is reminiscent of the Maximum Likelihood Amplitude Estimation (MLAE) approach to AE [11]. The main difference here is the classical post-processing, which provides far better results than MLAE. This was well-known for some time in the signal processing literature [21]. We used powers of two to determine scaling factors of our approach in Sec. IV.1, but note that there is a great deal of flexibility here depending on the use-case.
@@ III The AE algorithm
2:  Measure the state |$\ket{\phi_{n}}$ for $n\in D$ in the $Z$ basis and save the statistics $\hat{p_{0}}(n)$ ($\hat{p_{0}}(n)$ being the empirical probability of measuring |$\ket{0}$).
3:  Discretize the interval $[0,\pi/2]$ using $O(|D|)$ parts, learn the correct sign distribution and pick $\theta_{initial}$ in this discretization that is the most likely and use the signs of $\sin((2n+1)2\theta_{initial})$ as the initial guess, say $s^{0}=(s_{1},\dots,s_{|D|})$.
4:  for  $k$ in range($|D|$do
5:     Compute $S=\text{SlidingWindowSigns}(s^{0},k)$.
6:     for $s\in S$ do
7:        Form the “signal” vector $y_{n}=2\hat{p_{0}}(n)-1+is_{i}\sqrt{1-[2\hat{p_{0}}(n)-1]^{2}}$.
8:        Form the signal vector $\boldsymbol{y}\in\mathbb{C}^{M}$ by computing the signal values at the virtual locations corresponding to positive virtual positions.
9:        Use the ESPRIT (or any classical DOA) Algorithm 1 with $\boldsymbol{y}$ as input to extract the angle $\hat{\omega}$.
10:        Compute its negative log likelihood and save the signs and estimated angle $\hat{\omega}$.
11:     end for
12:  end for
13:  Return the angle $\hat{\omega}$ that minimized the objective function in the previous loop.
In classical signal processing terms, we obtain an estimate of the of the signal $y_{n}=e^{in4\theta+2i\theta}$ by placing a physical sensor at location $n$. Quantumly, we can obtain a estimate of the real-part of the signal $y_{n}$ by taking repeated measurements of |$\ket{\phi_{n}}$ in the $Z$ basis. As far as the DOA algorithms are concerned it does not matter how the measurements are obtained. However, as mentioned in Section II, we can’t quite distinguish between $y_{n}$ and $y_{n}^{*}$ from these measurement, since we can’t deduce the correct quadrant of $(2n+1)2\theta$: using the value of $\cos((2n+1)2\theta)$ we can only obtain the value of $|\sin((2n+1)2\theta)|$. It’s essential to obtain the correct sign of $\sin((2n+1)2\theta)$ so we can form the signal $y_{n}$.
III.1 Obtaining the correct signs
To obtain the correct sign of $\sin((2n+1)2\theta)$ for each $n$ (in your sequence of depths $D$), we need an extra subroutine that checks whether an assignment of signs is likely or not. For this, we need an objective function that assigns a score to each assignment of signs. Then we try many different assignments of signs (according to some heuristic) and we minimize the given objective function and use the signs that minimize the objective function to form the signal $y_{n}$. The algorithm that we propose for this purpose works in practice as our numerics will show, but can likely be further improved.
Given a sign vector and the estimated probabilities $\hat{p}_{0}(n)$ from Eq. (3) that we obtain by taking $N_{\textrm{shots}}(n)$ shots at each depth $n$, we can estimate $\theta_{est}$ using the ESPIRIT algorithm. From this estimate of $\theta_{est}$, we can compute an updated estimate $\hat{p}_{0}(n)$. For each of the different sign combinations, we minimize the negative log likelihood function
$\begin{split}-\log(\mathcal{L})&=-\sum_{n\in D}\bigg{\{}N_{0}(n)\log(\hat{p}_{0}(n))\\ &+[N_{0}(n)-N_{\textrm{shots}}(n)]\log(1-\hat{p}_{0}(n))\bigg{\}},\end{split}$ (13)
where $D$ is the set of depths at which we sample, $N_{0}(n)$ are the number of measurements that resulted in a 0, and $N_{\textrm{shots}}(n)$ are the total number of shots taken at depth $n$.
[…]
We still have to make a choice of starting vector of signs. The way we approach this is to first “learn” the signs for each angle $\theta\in\Theta\subset[0,\pi/2]$ where $\Theta$ is a discrete subset of angles. The starting vector of signs is then the one that minimizes the objective function in this set. The size of $\Theta$ can be taken to be $O(|D|)$ so that the classical complexity increases by another logarithmic factor of $O(\log(M))$.
III.2 Circuit Depth and Samples
The next step is to determine at what depths we take measurements. The DOA algorithms expects measurements on a uniform linear array of length $M$, that is measurements of $y_{n}$ for $n\in[M]$. One may obtain this by measuring |$\ket{\phi_{n}}$ at all depths $n\in[M]$, with a query complexity of $\sum_{n=1}^{M}n=O(M^{2})$. This does not provide us with the expected quantum scaling $O(1/M)$ of the error, but rather the scaling is $O(1/\sqrt{M})$ in this case. We avoid this problem by taking samples from a far smaller set of depths and use the virtual array concept to obtain (virtual) measurements at all depths $n\in[M]$.
@@ IV Numerics
Figure 3: Estimation error as a function of query complexity (top) and parallel complexity (bottom) for amplitude $a=0.5$. For csAE we simulated 500 runs with a sliding window value of length five giving a range of error rates (given by the vertical dots). The green triangles denote the 95-th percentile of the errors and the green fit line is used to compute the scaling factor. For chebAE we also simulated 500 runs, for a target $\varepsilon$ given by the results from csAE. The cloud of points from chebAE arise because of the iterative nature. The red and cyan stars and fit lines denote the average and worst case complexity respectively for chebAE for the 95-th percentile.
Table 1: Fitting constants for various confidence levels, along with the associated standard error obtained from the weighted-least-squares fit. The reported numbers are the average constant factors (resulting from the fit) on amplitudes sampled uniformly random in the interval $(0.1,0.9)$. Overall, csAE is comparable to the average case complexity of chebAE in terms of total query complexity, but tends to outperform in terms of worst-case complexity. In addition, csAE is highly parallelizable resulting in vast out-performance for parallel query complexity.
csAE $8.9\pm 0.9$ $0.27\pm 0.03$ $4.3\pm 0.1$ $0.26\pm 0.01$ $1.67\pm 0.05$ $0.102\pm 0.003$
chebAE (worst) $6.1\pm 0.2$ $5.87\pm 0.05$ $4.88\pm 0.05$ $4.3\pm 0.1$ $2.05\pm 0.02$ $1.79\pm 0.06$
chebAE (ave) $4.1\pm 0.2$ $3.9\pm 0.2$ $3.07\pm 0.11$ $2.95\pm 0.06$ $1.30\pm 0.01$ $1.19\pm 0.02$
@@
The query complexity is the number of times we use the operator $U$ (see  (1)) multiplied by two. Since the Grover operator $G$ consists of two calls to $U$, the query complexity will be the number of times we use $G$ plus the number of oracle calls at depth 0 divided by two. In chebAE [15] the number of oracle queries is defined as the number of times we apply the Grover operator $G$ ignoring the calls to $U$ at depth 0. This won’t affect constant factors in any meaningful way (we are penalizing our algorithm slightly). We compare the two approaches ability to achieve a target $\varepsilon$ and $\delta$ on two different metrics: 1) query complexity and 2) parallel query complexity. By parallel query complexity we mean the maximum depth oracle query that can be computed at compile time.
To get a better intuition of the performance of the algorithm, in Fig. 3, we provide results when estimating the amplitudes chosen uniformly at random over the range $a\in(0.1,0.9)$ for 500 trials. We use a physical array with parameters $N_{1}=N_{2}=\cdots N_{2q}=2$ as defined in Eq. (11) for $q\in\{3,4,\ldots,6\}$. For csAE, the input is the physical sampling array, denoting the fixed query depth, as well as the number of samples to take at each depth. This results in a deterministic schedule and yields the vertical colored dots representing the estimation errors of the algorithm from different runs. The green triangles represent the $\delta=0.95$ confidence level of those runs for different depth circuits. The green line is a fit to these 95% confidence levels.
@@
where $n_{\mathrm{max}}$ is the maximum query depth, $n$ is the current query depth, and $K$ is a constant that we chose to be $K=4$ for $\delta\leq 0.95$ and $K=8$ for $\delta=0.99$. It is possible that there exist better schedules $N_{\mathrm{shots}}(n)$. We note that it is also possible to have a non-deterministic schedule that utilizes information from the the top-two eigenvalues of the matrix $\boldsymbol{P}$ of Algorithm 1. When these eigenvalues are close to the remaining eigenvalues the signal and noise subspaces are not well separated. We believe that using this information could further improve the constant factor scaling, but at the cost of reducing parallelizability. We leave this to future work.
@@ IV.1 Fits
The performance of the algorithm depends on the value of the amplitude $a$ to be estimated. Since $a$ is an unknown parameter, we compare the performance of csAE for amplitudes sampled uniformly in the interval $(0.1,0.9)$. For each amplitude, we perform the same analysis shown in Fig. 3 and compute the estimation error $\varepsilon_{\delta}$ at a given confidence-level $\delta$. We compute the weighted-least squares fit using both average and worse-case query complexities for each amplitude to compute the scaling factor $C_{\textrm{opt}}$ and report the average number in Tab. 1. Here we provide a summary of the constant factor scaling comparing csAE to chebAE for different confidence levels and for both average and worst-case query complexity for chebAE.
Overall, we find that our algorithm under-performs chebAE for the average case, but matches the performance in the worst case to within statistical uncertainties. The greatest benefit from using our algorithm is its parallelizability, which greatly outperforms chebAE since that algorithm is inherently sequential.
@@
In addition to matching the performance of chebAE in terms of total oracle queries, the implementation of csAE is also simpler. As with the IQAE implementation [14], the csAE approach makes use of the standard Grover iterator and does not require Quantum Signal Processing (QSP). This implies that the circuit-level implementation is simpler than for chebAE, not requiring controlled phase rotations and circuit synthesis, which adds a logarithmic factor to the required fault-tolerant circuit depth [33]. Therefore, our approach yields the best of both worlds: simple circuit construction, together with state-of-the-art query complexity. As we discuss next, our approach also offers a highly parallel implementation, something not available to either chebAE or IQAE.
IV.2 Target Error Performance
Since the query depths and number of samples are pre-determined for csAE (which enables its parallelizability), it requires some trial and error if one wishes instead to achieve a desired target estimation error $\epsilon_{\delta}$ for a given confidence level $\delta$. The fits provide a useful parameter with which to quickly estimate the number of queries, but here we also provide concrete examples of the number of queries required to achieve a target error of $\epsilon_{\delta}=10^{-3}$ with confidence levels of $99\%$, $95\%$ and $68\%$, which are often used in practice when performing detailed resource estimation.
Results are presented in Tab. 2. As before, we estimate the performance by simulating the performance of the csAE algorithm for amplitudes sampled uniformly at random over the range $a=(0.1,0.9)$ using 500 Monte Carlo trials. These results provide a quick reference with which one can perform detailed resource estimation for their desired application.
Table 2: Sequential queries, maximum depth query, and simulated performance to achieve a target error rate of $\epsilon_{\delta}=10^{-3}$ for various confidence levels $\delta$. Details of the array settings and sampling strategy are given in Tabs. 5-5 in App. A.
Sequential Queries Maximum Depth Error
$\delta=0.99$ 8,777 256 $9.3\times 10^{-4}$
$\delta=0.95$ 4,488 256 $8.0\times 10^{-4}$
$\delta=0.68$ 1,560 128 $9.8\times 10^{-4}$
IV.3 Parallelizability
Figure 4: Required number of sequential applications of the circuit $U$ as we increase the number of quantum computers. This is for the amplitude $0.5$ to achieve the target error rate $\epsilon=10^{-3}$ at 95% confidence. The blue curve is for classical sampling where we simply apply the unitary $U$ once and sample. The black and green curves are for the worst and average case complexity of chebAE respectively. Here we assume at 2 QPUs chebAE can achieve maximum parallelizability, which is an optimistic assumption. Finally, the red curve is for csAE. Here we see the algorithm is fully parallel until it saturates. This occurs when the largest depth circuit is run on a single QPU for each shot. The remaining shorter depth circuits can be run somewhat sequentially on additional QPUs due to the exponential schedule.